# A Neural Difference-of-Entropies Estimator for Mutual Information

### v1

@@ 5 Conclusions
In this research, we proposed a new MI estimator which is based on the block autoregressive flow structure and the difference-of-entropies (DoE) estimator. Theoretically, our method converges to true mutual information as the number of samples increases and with large-enough neural network capacity. The accuracy of our estimator then depends on the ability of the block autoregressive flows in predicting the true posterior probability of items in the test set. A theoretical analysis is provided in the appendix. We discussed the connections and differences between our approach and other approaches including the lower bound approaches of MINE and SMILE and InfoNCE(CPC), and the classifier based approaches of CCMI and DEMI. We also demonstrate empirical advantages of our approach over the state of the art methods for estimating MI in synthetic data. Given its simplicity and promising performance, we believe that our method is a good candidate for use in research that optimizes MI. In future work, we aim to expand our experiments to additional data, including the image-like data of (Butakov et al., 2024) and the recently published benchmarks in (Lee & Rhee, 2024).
Despite its promising results, the proposed method has limitations in its current form. As a method that depends on the particular neural network architecture used to implement the flows, care needs to be taken to ensure the stability of the proposed estimator and its performance on smaller data sets. As seen in the experiments, the method performs particularly well in cases where the random quantities are based on Gaussian distribution. In future work we aim to explore the possibilities of using different classes of base distributions. This includes the possibility of dealing with discrete distributions, a situation that is handled well by critic-based methods Belghazi et al. (2018); van den Oord et al. (2018). Another direction involves evaluating our method in view of downstream applications that require the computation of mutual information and comparing its performance in these settings with other generative approaches that were recently introduced (Franzese et al., 2024; Duong & Nguyen, 2023; Butakov et al., 2024). In particular, in light of recent work Kong et al. (2023); Franzese et al. (2024), it would be interesting to explore multimodal examples, like the MI between image data and text embeddings.
@@ Appendix A Real NVP
An alternative architecture for implementing our approach is based on Real NVP. Real NVP, proposed by Dinh et al. (2016), is a class of normalizing flows constructed using simple and flexible bijections with efficient computation of the Jacobian determinant. Each transformation in Real NVP is referred to as an affine coupling layer. Given a $d$-dimensional input $x$ and a partition point $d_{m}<d$, the output $y$ of an affine coupling layer is defined by the following equations:
$y_{1:d_{m}}=x_{1:d_{m}},\ y_{d_{m}+1:d}=x_{d_{m}+1:d}\odot\exp(s(x_{1:d_{m}}))+t(x_{1:d_{m}}),$
@@
$\begin{bmatrix}\mathbb{I}_{d_{m}}&0\\ \frac{\partial y_{d_{m}+1:d}}{\partial x_{1:d_{m}}^{T}}&\textnormal{diag}(\exp(s(x_{1:d_{m}})))\end{bmatrix},$
where $\textnormal{diag}(\exp(s(x_{1:d_{m}})))$ is a diagonal matrix whose diagonal elements correspond to the vector $\exp(s(x_{1:d_{m}}))$. As the Jacobian is triangular, its determinant can be efficiently computed as $\exp(\sum_{j}s(x_{1:d_{m}})_{j})$. Additionally, since the computation does not depend on the Jacobian of $s$ and $t$, arbitrarily complex functions can be used for $s$ and $t$. A common choice is deep convolutional neural networks with more features in the hidden layers than in the input and output layers.
@@ Appendix B Theoretical results
with $x\in\mathbb{R}^{n}$, $y\in\mathbb{R}^{n}$. The corresponding flows are implemented by block-triangular maps. We call $T\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ triangular, if the following conditions are satisfied for $i\in\{1,\dots,n\}$:
@@
    For any $x_{1:i-1}=(x_{1},\dots,x_{i-1})$, $T_{i}(x_{1:i-1},x_{i})$ is monotonically increasing as a function of $x_{i}$.
It is easy to see that triangular maps are invertible. It is a well-known result that normalizing flows can be implemented as triangular maps (Bogachev et al., 2007). In our setting, we are interested in block-triangular maps of a specific form. We call a map $T\colon\mathbb{R}^{2n}\to\mathbb{R}^{2n}$ block triangular, if $T=(T_{1}(y),T_{2}(y,x))$, where each $T_{i}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $\det\mathrm{d}T_{1}(y)>0$, as well as $\det\mathrm{d}_{y}T_{2}(y,x)>0$ for each $y$ and $x$. While in our case we restrict to the case where $x$ and $y$ have the same dimension, everything that follows generalizes to different dimensions. Recall that if $T\colon\mathbb{R}^{2n}\to\mathbb{R}^{2n}$ is a map and $\mu$ is a probability measure on $\mathbb{R}^{2n}$, then the push-forward measure is given by
@@
We note that the existence of $T_{1}$ and $T_{2}$ as in the theorem follows from Bogachev et al. (2007) (they may even be chosen to be triangular maps).
@@
Let $\overline{\mu}=U_{*}\mu$. By Bogachev et al. (2007), we know that there exists a triangular transformation $V=(T_{1},\overline{T})\colon\mathbb{R}^{2n}\to\mathbb{R}^{2n}$, with $\overline{T}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $V_{*}\overline{\mu}=\nu$. It follows that
@@
Let $f^{x}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a normalizing flow for the marginal density $p_{X}(x)$. The there exists a block-triangular map $f=(f_{1},f_{2})$ that is a normalizing flow for $p(x,y)$, where $f_{2}(y,x)=\overline{f}(y,f^{x}(x))$ for a map $\overline{f}\colon\mathbb{R}^{2n}\to\mathbb{R}^{n}$, and such that
$p(x|y)=q(\overline{f}(y,f^{x}(x)))\cdot|\det\mathrm{d}_{x}f_{2}(y,x)|,$
@@
The idea is to implement such a flow using neural networks, in such a way that by deactivating a certain part of the neural network for $f_{2}(y,x)=\overline{f}(y,f^{x}(x))$, we get the flow $f^{x}$. In analogy to Block Neural Autoregressive Flows, we now consider
@@
with ${B}^{k}_{ij}\in\mathbb{R}^{m_{k}\times m_{k-1}}$, $g=\exp$ and $m_{0}=d$. For $k=\ell$ we omit the activation function and set $m_{k}=d$. The following basic result is shown along the lines of De Cao et al. (2019).
@@
Moreover, if $\sigma$ is strictly increasing, then $\det\mathrm{d}_{{y}}f_{1}({y})>0$ and $\det\mathrm{d}_{{x}}f_{2}({y},{x})>0$ for all $({y},{x})$.
@@
we can obtain the entropy associated to the marginal density, $H(X)$, by “deactivating” the weights that operate on $y$. Corollary B.2 suggests that given enough expressive power of our neural network architecture, we can train the network to both approximate $H(X|Y)$ and $H(X)$ by first training the marginal density $f^{x}$ and then training for $(f_{1},\overline{f})$ on samples $(y_{i},f^{x}(x_{i}))$. Along the lines of the appendix in De Cao et al. (2019), one can show that any conditional density can be approximated by a block-triangular flow as described.
@@ C.1 MI Estimation with Nonlinear Transformations
The first experiment we conducted are focused on estimating MI with samples that are generated with nonlinearly transformed Gaussians. Here we consider asinh and wiggly transformations that are provided in (Czyż et al., 2023). The different MI estimators were trained on the $20$-dimensional datasets of the varying sample size ($32$K, $64$K and $128$K). All the other settings remained the same.
@@ C.3 MI Estimation on Correlated Uniforms and Student’s t Distributions
It is well-known that non-Gaussian distributions, especially distributions with long tails, remains challenging for MI estimation for many reasons. In this experiment, we sampled from two random variables $X$ and $Y$ of correlated Uniform and Student’s t distributions, for which the MI can be exactly obtained from their known correlation. The construction of Student’s t distribution follows the idea of Czyż et al. (2023) where the Gaussians are taken from the first experiment. Note that, $I(X,Y)>0$ for the generated Student’s t distribution even for independent Gaussians $X,Y$ in this example. We trained the different estimators on the training set of 128K samples and tested them on another independently generated 10240 samples to obtain MI estimates. All the other settings remained the same.
@@ C.4 Comparison between NDoE and BNAF
Results. As shown in the plots, NDoE, BNAF demonstrated better convergence behavior with relatively small error after just 10 epochs of training, whereas BNAF required between 20 to 60 epochs to achieve competitive results. This performance gap is particularly pronounced when the number of training samples is smaller. Another noteworthy observation is that NDoE, BNAF empirically behaved as an upper-bound estimator, consistently producing estimates greater than zero, which aligns well with the fundamental properties of mutual information (MI). In contrast, BNAF does not exhibit this property, which is often regarded as a key limitation of generative methods. We attribute this phenomenon to the Correlation Boosting Effect proposed by Gao et al. (2017), though we do not provide a rigorous proof at this stage.
@@ C.5 Long-run Training Behavior of NDoE
We noticed from the last three experiments that DEMI underestimates the MI when the random variables are highly dependent. The underestimation is not alleviated by the increased number of training epochs. At the same time, our method shows the reduction of bias with the repetition of training. Thus, we proposed an assumption that the discriminative methods diverge to the true MI for high mutual dependence, while generative models have a good convergence property with sufficient training samples, with the precondition of enough expressive power of the neural network. In other words, there is a systematic bias in the discriminative methods which are positively correlated to mutual dependence. To verify this experimentally, we conducted a long-run training behavior experiment using similar settings of Liao et al. (2020). Here we only verify the long-run training behavior for NDoE, BNAF. All the estimators were trained on the 20-dimensional Gaussian and Cubed Gaussian case for 100000 training steps with batch size 128. Samples were drawn directly from the generating distributions. We did this for four ground-truth MI values of $0.1,10,20$, and $30$.
Results. The results are shown in Figure 13 and Figure 14. Among all the methods, our method shows the best convergence behaviour for all MI with the increase of training epochs, and the variances remains relatively small DEMI method is competitive when true MI is close to 00. However, it underestimates MI vastly for large true MI. We also noticed that SMILE with the parameter $\tau=1,5$ tends to overestimate MI even though they are based on a lower bound of it. McAllester & Stratos (2018) suggests that the reasoning behind this could be the sensitivity of the estimate of $-\textnormal{ln}\ \mathbb{E}[e^{f(x,y)}]$ to outliers. However, the parameter $\tau$ actually clips the term $e^{f(x,y)}$ to the interval $[e^{-\tau},e^{\tau}]$, which removes outliers from the neural network outputs.
@@ C.6 Asymmetry test and Varying Batchsize
It is obvious to see that our proposed method is asymmetric, i.e. estimating the MI by using $I(X,Y)=H(X)-H(X,Y)$ and $I(X,Y)=H(Y)-H(Y,X)$ could obtain different results. We did an extra experiment on the above distribution to show that the difference is minor. The results are shown in Figure 15. Since our method tends to underestimate the true MI in the experiments, choosing the larger estimation will lend to less bias in most cases. Another experiment focuses on varying batchsize. It is believed that the poor performance of discriminative methods on high MI estimation are dependent on the batchsize. We choose the batchsize of $64,128,256$ and $512$ to see the performance of our method and compare them with the baselines. The results are shown in Figure 16.
@@ Appendix D Self-consistency
In applications with real data, obtaining the ground truth MI is challenging or not possible. However, as suggested by Song & Ermon (2019), one can still test whether a MI estimator satisfies some of the fundamental properties of MI: $I(X,Y)=0$ if $X$ and $Y$ are independent, the data processing inequality is satisfied (that is, transforming $X$ and $Y$ should not increase the MI), and additivity.
Following Song & Ermon (2019), we conducted self-consistency tests on high-dimensional images (MNIST) under three settings, where obtaining the ground truth MI is challenging. These settings involve processing images $X$ and $Y$ in different ways to assess the performance of various methods: DEMI, InfoNCE, SMILE, NDoE, BNAF and BNAF, where NDoE, BNAF and BNAF applies autoencoders (AE) for dimensionality reduction. DoE is not included as it is considered for certain failure in the experiments before. The three settings include:
@@
    Additivity. $X$ corresponds to two independent images, and $Y$ includes the top t rows of both. The assessment focuses on the estimated MI ratio between $[X_{1},X_{2}]$ and $[Y_{1},Y_{2}]$ relative to the true MI between $X$ and $Y$.
@@
For the data-processing test, we set $t_{2}=t_{1}-3$. Ideally, the estimator should satisfy $\hat{I}([X,X];[Y,h(Y)])/\hat{I}(X,Y)\approx 1$. This is because additional processing should not result in an increase in information. All methods performs relatively well except for NDoE, BNAF and BNAF. This is possibly due to limited capacity of AE.
In the additivity setting, the estimator should ideally double its value compared to the baseline with the same t, i.e. $\hat{I}([X_{1},X_{2}];[Y_{1},Y_{2}])/\hat{I}(X,Y)\approx 2$. Discriminative approaches did not perform well in this case, except when t was very small. As t increased, this ratio converged to 1, possibly due to initialization and saturation of the training objective. However, NDoE, BNAF performed well on this test except when t is small ($t=0,3$). Compared with the results from Song & Ermon (2019), it is promising to see improving performance by using VAE instead.
Refer to captionRows UsedRatioMNIST (Baseline)
Refer to captionRows UsedRatioMNIST (Data Processing)
Refer to captionRows usedRatioMNIST (Additivity)
@@
Figure 17: Evaluation on high-dimensional images (MNIST) under three settings. From top to bottom: Evaluation of $\hat{I}(X;Y)/\hat{I}(X;X)$; Evaluation of $\hat{I}([X,X];[Y,h(Y)]/\hat{I}(X;Y)$, where the ideal value is 1; Evaluation of $\hat{I}([X_{1},X_{2}];[Y_{1},Y_{2}]/\hat{I}(X;Y)$, where the ideal value is 1. $X$ is an image, $Y$ contains the top $t$ rows of $X$ and $h(Y)$ contains the top $(t-3)$ rows of $X$.

### v3

@@ 5 Conclusions
In this research, we proposed a new MI estimator which is based on the block autoregressive flow structure and the difference-of-entropies (DoE) estimator. Theoretically, our method converges to true mutual information as the number of samples increases and with large-enough neural network capacity. The accuracy of our estimator then depends on the ability of the block autoregressive flows in predicting the true posterior probability of items in the test set. A theoretical analysis is provided in Section B of supplementary materials [26]. We discussed the connections and differences between our approach and other approaches, including the lower bound approaches of MINE and SMILE, InfoNCE (CPC), and the classifier-based approaches of CCMI and DEMI. We also demonstrate empirical advantages of our approach over the state-of-the-art methods for estimating MI in synthetic data. Given its simplicity and promising performance, we believe that our method is a good candidate for use in research that optimizes MI. In future work, we aim to expand our experiments to additional data, including the image-like data of Butakov et al. [4] and the recently published benchmarks in Lee and Rhee [19].
Despite its promising results, the proposed method has limitations in its current form. As a method that depends on the particular neural network architecture used to implement the flows, care needs to be taken to ensure the stability of the proposed estimator and its performance on smaller data sets. As seen in the experiments, the method performs particularly well in cases where the random quantities are based on Gaussian distribution. In future work we aim to explore the possibilities of using different classes of base distributions. This includes the possibility of dealing with discrete distributions, a situation that is handled well by critic-based methods [2, 33]. Another direction involves evaluating our method in view of downstream applications that require the computation of mutual information and comparing its performance in these settings with other generative approaches that were recently introduced [11, 10, 4]. In particular, in light of recent work [17, 11], it would be interesting to explore multimodal examples, such as the MI between image data and text embeddings.
{ack}
The first author was funded by a China Scholarship Council scholarship.
@@ Appendix A Real NVP
An alternative architecture for implementing our approach is based on Real NVP. Real NVP, proposed by Dinh et al. [9], is a class of normalizing flows constructed using simple and flexible bijections with efficient computation of the Jacobian determinant. Each transformation in Real NVP is referred to as an affine coupling layer. Given a $d$-dimensional input $x$ and a partition point $d_{m}<d$, the output $y$ of an affine coupling layer is defined by the following equations:
$y_{1\mathrel{\mathop{\ordinarycolon}}d_{m}}=x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}},\ y_{d_{m}+1\mathrel{\mathop{\ordinarycolon}}d}=x_{d_{m}+1\mathrel{\mathop{\ordinarycolon}}d}\odot\exp(s(x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}}))+t(x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}}),$
@@
$\begin{bmatrix}\mathbb{I}_{d_{m}}&0\\ \frac{\partial y_{d_{m}+1\mathrel{\mathop{\ordinarycolon}}d}}{\partial x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}}^{T}}&\textnormal{diag}(\exp(s(x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}})))\end{bmatrix}$,
where $\textnormal{diag}(\exp(s(x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}})))$ is a diagonal matrix whose diagonal elements correspond to the vector $\exp(s(x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}}))$. As the Jacobian is triangular, its determinant can be efficiently computed as $\exp(\sum_{j}s(x_{1\mathrel{\mathop{\ordinarycolon}}d_{m}})_{j})$. Additionally, since the computation does not depend on the Jacobian of $s$ and $t$, arbitrarily complex functions can be used for $s$ and $t$. A common choice is deep convolutional neural networks with more features in the hidden layers than in the input and output layers.
@@ Appendix B Theoretical results
with $x\in\mathbb{R}^{n}$, $y\in\mathbb{R}^{n}$. The corresponding flows are implemented by block-triangular maps. We call $T\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ triangular if the following conditions are satisfied for $i\in\{1,\dots,n\}$:
@@
    For any $x_{1\mathrel{\mathop{\ordinarycolon}}i-1}=(x_{1},\dots,x_{i-1})$, $T_{i}(x_{1\mathrel{\mathop{\ordinarycolon}}i-1},x_{i})$ is monotonically increasing as a function of $x_{i}$.
It is easy to see that triangular maps are invertible. It is a well-known result that normalizing flows can be implemented as triangular maps [3]. In our setting, we are interested in block-triangular maps of a specific form. We call a map $T\colon\mathbb{R}^{2n}\to\mathbb{R}^{2n}$ block triangular, if $T=(T_{1}(y),T_{2}(y,x))$, where each $T_{i}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $\det\mathrm{d}T_{1}(y)>0$, as well as $\det\mathrm{d}_{y}T_{2}(y,x)>0$ for each $y$ and $x$. While in our case we restrict to the case where $x$ and $y$ have the same dimension, everything that follows generalizes to different dimensions. Recall that if $T\colon\mathbb{R}^{2n}\to\mathbb{R}^{2n}$ is a map and $\mu$ is a probability measure on $\mathbb{R}^{2n}$, then the push-forward measure is given by
@@
We note that the existence of $T_{1}$ and $T_{2}$ as in the theorem follows from Bogachev et al. [3] (they may even be chosen to be triangular maps).
@@
Let $\overline{\mu}=U_{*}\mu$. By Bogachev et al. [3], we know that there exists a triangular transformation $V=(T_{1},\overline{T})\colon\mathbb{R}^{2n}\to\mathbb{R}^{2n}$, with $\overline{T}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $V_{*}\overline{\mu}=\nu$. It follows that
@@
Let $f_{(x)}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a normalizing flow for the marginal density $p_{X}(x)$. There exists a block-triangular map $f=(f_{1},f_{2})$ that is a normalizing flow for $p(x,y)$, where $f_{2}(y,x)=\overline{f}(y,f_{(x)}(x))$ for a map $\overline{f}\colon\mathbb{R}^{2n}\to\mathbb{R}^{n}$, and such that
$p(x|y)=q(\overline{f}(y,f_{(x)}(x)))\cdot|\det\mathrm{d}_{x}f_{2}(y,x)|,$
@@
The idea is to implement such a flow using neural networks, in such a way that by deactivating a certain part of the neural network for $f_{2}(y,x)=\overline{f}(y,f_{(x)}(x))$, we get the flow $f_{(x)}$. In analogy to Block Neural Autoregressive Flows, we now consider
@@
with ${B}^{k}_{ij}\in\mathbb{R}^{m_{k}\times m_{k-1}}$, $g=\exp$ and $m_{0}=d$. For $k=\ell$ we omit the activation function and set $m_{k}=d$. The following basic result is shown along the lines of De Cao et al. [8].
@@
Moreover, if $\sigma$ is strictly increasing, then $\det\mathrm{d}_{{y}}f_{1}({y})>0$ and $\det\mathrm{d}_{{x}}f_{2}({y},{x})>0$ for all (y,x)({y},{x}).
@@
we can obtain the entropy associated to the marginal density, $H(X)$, by “deactivating” the weights that operate on $y$. Corollary B.2 suggests that given enough expressive power of our neural network architecture, we can train the network to both approximate $H(X|Y)$ and $H(X)$ by first training the marginal density $f_{(x)}$ and then training for $(f_{1},\overline{f})$ on samples $(y_{i},f_{(x)}(x_{i}))$. Along the lines of the appendix in De Cao et al. [8], one can show that any conditional density can be approximated by a block-triangular flow as described.
@@ C.1 MI Estimation with Nonlinear Transformations
The first experiment we conducted are focused on estimating MI with samples that are generated with nonlinearly transformed Gaussians. Here we consider asinh and wiggly transformations that are provided in [7]. The different MI estimators were trained on the $20$-dimensional datasets of the varying sample size (3232K, 6464K and 128128K). All the other settings remained the same.
@@ C.3 MI Estimation on Correlated Uniforms and Student’s t Distributions
It is well-known that non-Gaussian distributions, especially distributions with long tails, remains challenging for MI estimation for many reasons. In this experiment, we sampled from two random variables $X$ and $Y$ of correlated Uniform and Student’s t distributions, for which the MI can be exactly obtained from their known correlation. The construction of Student’s t distribution follows the idea of Czyż et al. [7] where the Gaussians are taken from the first experiment. Note that, $I(X,Y)>0$ for the generated Student’s t distribution even for independent Gaussians $X,Y$ in this example. We trained the different estimators on the training set of 128K samples and tested them on another independently generated 10240 samples to obtain MI estimates. All the other settings remained the same.
@@ C.4 Comparison between NDoE and BNAF
Results. As shown in the plots, NDoE, BNAF demonstrated better convergence behavior with relatively small error after just 10 epochs of training, whereas BNAF required between 20 to 60 epochs to achieve competitive results. This performance gap is particularly pronounced when the number of training samples is smaller. Another noteworthy observation is that NDoE, BNAF empirically behaved as an upper-bound estimator, consistently producing estimates greater than zero, which aligns well with the fundamental properties of mutual information (MI). In contrast, BNAF does not exhibit this property, which is often regarded as a key limitation of generative methods. We attribute this phenomenon to the Correlation Boosting Effect proposed by Gao et al. [12], though we do not provide a rigorous proof at this stage.
@@ C.5 Long-run Training Behavior of NDoE
We noticed from the last three experiments that DEMI underestimates the MI when the random variables are highly dependent. The underestimation is not alleviated by the increased number of training epochs. At the same time, our method shows the reduction of bias with the repetition of training. Thus, we proposed an assumption that the discriminative methods diverge to the true MI for high mutual dependence, while generative models have a good convergence property with sufficient training samples, with the precondition of enough expressive power of the neural network. In other words, there is a systematic bias in the discriminative methods which are positively correlated to mutual dependence. To verify this experimentally, we conducted a long-run training behavior experiment using similar settings of Liao et al. [20]. Here we only verify the long-run training behavior for NDoE, BNAF. All the estimators were trained on the 20-dimensional Gaussian and Cubed Gaussian case for 100000 training steps with a batch size of $128$. Samples were drawn directly from the generating distributions. We did this for four ground-truth MI values of $0.1,10,20$, and $30$.
Results. The results are shown in Figure 13 and Figure 14. Among all the methods, our method shows the best convergence behaviour for all MI with the increase of training epochs, and the variances remains relatively small DEMI method is competitive when true MI is close to 0. However, it underestimates MI vastly for large true MI. We also noticed that SMILE with the parameter $\tau=1,5$ tends to overestimate MI even though they are based on a lower bound of it. McAllester and Stratos [23] suggests that the reasoning behind this could be the sensitivity of the estimate of $\textnormal{ln}\ \mathbb{E}[e^{f(x,y)}]$ to outliers. However, the parameter $\tau$ actually clips the term $e^{f(x,y)}$ to the interval $e^{-\tau},e^{\tau}]$, which removes outliers from the neural network outputs.
@@ C.6 Asymmetry test and Varying Batchsize
It is obvious to see that our proposed method is asymmetric, i.e. estimating the MI by using I(X,Y)=H(X)H(X,Y)I(X,Y)=H(X)-H(X,Y) and I(X,Y)=H(Y)H(Y,X)I(X,Y)=H(Y)-H(Y,X) could obtain different results. We did an extra experiment on the above distribution to show that the difference is minor. The results are shown in Figure 15. Since our method tends to underestimate the true MI in the experiments, choosing the larger estimation will lead to less bias in most cases. Another experiment focuses on varying batchsize. It is believed that the poor performance of discriminative methods on high MI estimation is dependent on the batch size. We choose the batchsize of $64,128,256$ and $512$ to see the performance of our method and compare them with the baselines. The results are shown in Figure 16.
@@ Appendix D Self-consistency
In applications with real data, obtaining the ground truth MI is challenging or not possible. However, as suggested by Song and Ermon [31], one can still test whether a MI estimator satisfies some of the fundamental properties of MI: $I(X,Y)=0$ if $X$ and $Y$ are independent, the data processing inequality is satisfied (that is, transforming $X$ and $Y$ should not increase the MI), and additivity.
Following Song and Ermon [31], we conducted self-consistency tests on high-dimensional images (MNIST) under three settings, where obtaining the ground truth MI is challenging. These settings involve processing images $X$ and $Y$ in different ways to assess the performance of various methods: DEMI, InfoNCE, SMILE, NDoE, BNAF and BNAF, where NDoE, BNAF and BNAF apply autoencoders (AE) for dimensionality reduction. DoE is not included as it is considered for certain failure in the experiments before. The three settings include:
@@
    Additivity. $X$ corresponds to two independent images, and $Y$ includes the top t rows of both. The assessment focuses on the estimated MI ratio between $X_{1},X_{2}]$ and $Y_{1},Y_{2}]$ relative to the true MI between $X$ and $Y$.
@@
For the data-processing test, we set $t_{2}=t_{1}-3$. Ideally, the estimator should satisfy $\hat{I}([X,X];[Y,h(Y)])/\hat{I}(X,Y)\approx 1$. This is because additional processing should not result in an increase in information. All methods perform relatively well except for NDoE, BNAF and BNAF. This is possibly due to the limited capacity of AE.
In the additivity setting, the estimator should ideally double its value compared to the baseline with the same t, i.e. $\hat{I}([X_{1},X_{2}];[Y_{1},Y_{2}])/\hat{I}(X,Y)\approx 2$. Discriminative approaches did not perform well in this case, except when t was very small. As t increased, this ratio converged to 1, possibly due to initialization and saturation of the training objective. However, NDoE, BNAF performed well on this test except when t is small ($t=0,3$). Compared with the results from Song and Ermon [31], it is promising to see improved performance by using VAE instead.
@@
Figure 17: Evaluation on high-dimensional images (MNIST) under three settings. From left to right: Evaluation of $\hat{I}(X;Y)/\hat{I}(X;X)$; Evaluation of $\hat{I}([X,X];[Y,h(Y)]/\hat{I}(X;Y)$, where the ideal value is 1; Evaluation of $\hat{I}([X_{1},X_{2}];[Y_{1},Y_{2}]/\hat{I}(X;Y)$, where the ideal value is 1. $X$ is an image, $Y$ contains the top $t$ rows of $X$ and $h(Y)$ contains the top (t3)(t-3) rows of $X$.